# A flexible class of latent variable models for the analysis of antibody response data

### v2

@@ 6 Application to malaria serology
We present a re-analysis of malaria serology data from Bousema2013; Bousema2016, collected in 2011 from communities in Rachuonyo South District, western Kenyan highlands. Finger-prick blood samples were collected on filter paper and used to detect total immunoglobulin G (IgG) antibodies against two Plasmodium falciparum blood-stage antigens: apical membrane antigen 1 (AMA1) and merozoite surface protein 1 (MSP1). The outcome variable in our analysis is the optical density (OD) measured by ELISA.
@@ 6.1 Modelling apical membrane antigen 1 (AMA1) concentrations
Table 4 summarises the parameter estimates and uncertainty from the parametric bootstrap procedure for the joint model fitted across the full age range. The estimated change point for the transition from the mechanistic formulation to the regression-based formulation is $\hat{\tau}=20.8$ years. This suggests that the mechanistic framework remains appropriate up to approximately age 20. This finding demonstrates that rather than imposing an a priori age restriction, as is sometimes done in serology studies (e.g. yman2016), our model provides an empirical, data-driven approach to determining the appropriate age range for mechanistic modelling.
@@ 6.2 Modelling merozoite surface protein 1 (MSP1) concentrations
Figure 8 displays the expected mean of $Y$ as a function of age for both MSP1 and AMA1. AMA1 rises sharply in early childhood but then declines after about 20 years, which may reflect the higher individual-level variability in AMA1 antibody dynamics, with rapid boosting and waning following each infection (akpogheneta2008; wipasa2010; ondigo2014). The peak around 20 years may reflect the higher prevalence of asymptomatic infections with detectable parasitemia in this age range. At this stage, individuals have typically acquired clinical immunity (protection against disease symptoms) but have not yet fully developed anti-parasite immunity, sustaining infections with parasite densities high enough to continue boosting antibody levels. In contrast, as anti-parasite immunity develops in adults, lower parasite densities result in reduced antigenic stimulation and gradual antibody decline.
In contrast, MSP1 shows a steady increase across the full age range with no decline. This is consistent with MSP1 eliciting durable antibody responses (akpogheneta2008; wipasa2010). The change-point at $\zeta=11.6$ years marks an acceleration in the rate of increase. While the precise mechanism underlying this shift is unclear, plausible explanations include cumulative exposure effects associated with school attendance, immune maturation during early adolescence, or cohort effects reflecting temporal changes in malaria transmission intensity. Distinguishing among these possibilities would require longitudinal data or additional historical information on transmission patterns and intervention coverage.
@@ 7 Discussion
We have introduced a latent variable modelling framework that offers four key methodological contributions. First, it provides a principled way to model the full distribution of antibody concentrations without dichotomization, which remains the dominant practice in serological analyses (corran2007; Drakeley2005; Arnold2014; cox2022) despite well-established information loss (royston2006; kyomuhangi2021). By modelling antibody levels directly, the framework provides a more refined characterisation of how immune responses vary and evolve with age. Second, it offers a flexible approach for incorporating age-dependent structure through either mechanistic specifications that encode epidemiological assumptions about exposure and immune response dynamics, or data-driven specifications that adaptively capture empirical patterns without imposing strong structural constraints. Third, by explicitly representing immune activation as a continuum, the framework more naturally accommodates complex population heterogeneity in immunological responses. Fourth, we introduce an $L_{2}$ histogram-based estimator that provides a computationally efficient alternative to maximum likelihood estimation, enabling practical implementation of the framework even in computationally demanding contexts with large sample sizes.
@@
An important consideration for applying our framework concerns vaccination status. For diseases conferring durable immunity following vaccination or natural infection - e.g. yellow fever, where a single vaccine dose provides lifelong protection (Staples2015) - binary serological classification may suffice for most purposes, and continuous antibody modeling may offer limited added value. However, for diseases characterized by partial or waning immunity, such as malaria, continuous antibody measurements provide information about immune activation dynamics that is lost when data are dichotomized. The recent introduction of malaria vaccines (RTS,S/AS01 and R21/Matrix-M) (RTS2015; Datoo2021) presents new challenges for serological surveillance that affect any analytical approach, not just our modeling framework. Because these vaccines target the pre-erythrocytic circumsporozoite protein (CSP) rather than the blood-stage antigens we analyze (AMA1, MSP1), vaccination does not directly boost these antibody responses. However, reduced transmission resulting from vaccination programs could alter natural boosting patterns, thereby affecting the age-dependent dynamics we model. Our framework could be extended to incorporate vaccination status as a covariate affecting the latent immune state $T$, or to jointly model multiple antigens distinguishing vaccine-induced from infection-induced response extensions that will become increasingly important as malaria vaccination expands across sub-Saharan Africa.
@@
Future research will focus on extending the current modelling framework to address a wider range of questions in sero epidemiological research, including the development of joint models for multiple antibody responses. One natural direction is to replace the single latent quantity with a multivariate latent structure and to combine this with an appropriate multivariate Beta distribution, for example using the construction proposed by arnold2011. This would allow the model to capture dependencies between antibody responses to different antigens and provide deeper insight into how immune markers coevolve following exposure. The same idea extends naturally beyond serology, since the latent variable formulation is directly relevant to high dimensional omics settings such as proteomics and transcriptomics, where disease states are typically treated as discrete categories, for example healthy versus diseased or sepsis molecular endotypes. Recent work using neural network models for disease classification from multi omics data (hartman2023) and large transcriptomic and proteomic studies of sepsis subtypes (wong2012; reyes2020) all adopt this discrete label perspective, despite the underlying host response being biologically continuous rather than strictly partitioned. Within our framework this continuum can be represented by a latent variable $T$, with each biomarker, here each transcript or protein, modelled as an observed outcome $Y_{j}$ whose conditional distribution depends smoothly on $T$ through marker specific mean and variance functions that describe how expression evolves across the response axis. While a single latent variable provides a parsimonious representation of the dominant gradient of biological variation, many omics settings involve several distinct processes, for example inflammation, immune suppression, and endothelial injury, which may jointly determine the multivariate profile. In such cases it may be necessary to adopt a multivariate latent representation and to define the conditional distribution of $Y_{j}$ as a function of this collection of latent quantities, allowing the framework to capture more complex patterns and multiple dimensions of biological variation that cannot be reduced to a single continuum.
@@ Acknowledgements
Let $f_{\vartheta,n}:=\;f_{n}(\cdot;\vartheta)$. We apply Corollary 3.2.3 of vandervaartwellner1996. This corollary establishes consistency provided that: (i) the empirical criterion converges uniformly to the population criterion, i.e., $\sup_{\vartheta\in\vartheta}|\mathbb{M}_{n}(\vartheta)-\mathbb{M}(\vartheta)|\xrightarrow{p}0$. and that (ii) the population maximum is well-separated. Assumption A2. directly gives condition (ii). Thus, it remains only to show condition (i).
We prove this condition in two step. We first establish that the unparameterized histogram $\widehat{f}_{n}$ converges to the true density $f_{0}$ in the $L_{2}$ norm. freedman1981histogram provide the exact mean integrated squared error (MISE) decomposition in their Proposition (1.10):
@@
where $f_{0,n}$ is the theoretical discretized density (the expectation of $\widehat{f}_{n}$). The first term is bounded by $(nh_{n})^{-1}$ and under Assumption A4., this term vanishes. The second term corresponds to the approximation error of the binned density; as shown in Eq. (2.5) of freedman1981histogram, this vanishes as $h_{n}\to 0$ for any $f_{0}\in L_{2}$. Therefore, $\mathbb{E}[\|\widehat{f}_{n}-f_{0}\|_{2}^{2}]\to 0$, and by Markov’s inequality:
@@
Again, following Assumption A3. and (16), the terms in parentheses are bounded uniformly and bounded in probability (note that if two terms tend to zero in probability, so does their product). freedman1981histogram show in Proposition (2.7) that when the density is absolutely continuous with square-integrable derivative, then $\|f_{\vartheta,n}-f(\cdot;\vartheta)\|^{2}_{2}$ is $O(h_{n}^{2})$. Specifically, for small $h_{n}$, the term is bounded by a constant times $h_{n}^{2}\|f^{\prime}(\cdot;\vartheta)\|_{2}^{2}$. Thus under Assumption A5., $\sup_{\vartheta\in\vartheta}\|f^{\prime}(\cdot;\vartheta)\|_{2}<\infty$:

### v3

@@ 6 Application to malaria serology
We present a re-analysis of malaria serology data from Bousema et al. (2013, 2016), collected in 2011 from communities in Rachuonyo South District, western Kenyan highlands. Finger-prick blood samples were collected on filter paper and used to detect total immunoglobulin G (IgG) antibodies against two Plasmodium falciparum blood-stage antigens: apical membrane antigen 1 (AMA1) and merozoite surface protein 1 (MSP1). The outcome variable in our analysis is the optical density (OD) measured by ELISA.
@@ 6.1 Modelling apical membrane antigen 1 (AMA1) concentrations
Table 7 summarises the parameter estimates and uncertainty from the parametric bootstrap procedure for the joint model fitted across the full age range. The estimated change point for the transition from the mechanistic formulation to the regression-based formulation is $\hat{\tau}=20.8$ years. This suggests that the mechanistic framework remains appropriate up to approximately age 20. This finding demonstrates that rather than imposing an a priori age restriction, as is sometimes done in serology studies (e.g. Yman et al. (2016)), our model provides an empirical, data-driven approach to determining the appropriate age range for mechanistic modelling.
@@ 6.2 Modelling merozoite surface protein 1 (MSP1) concentrations
Figure 8 displays the expected mean of $Y$ as a function of age for both MSP1 and AMA1. AMA1 rises sharply in early childhood but then declines after about 20 years, which may reflect the higher individual-level variability in AMA1 antibody dynamics, with rapid boosting and waning following each infection (Akpogheneta et al., 2008; Wipasa et al., 2010; Ondigo et al., 2014). The peak around 20 years may reflect the higher prevalence of asymptomatic infections with detectable parasitemia in this age range. At this stage, individuals have typically acquired clinical immunity (protection against disease symptoms) but have not yet fully developed anti-parasite immunity, sustaining infections with parasite densities high enough to continue boosting antibody levels. In contrast, as anti-parasite immunity develops in adults, lower parasite densities result in reduced antigenic stimulation and gradual antibody decline.
In contrast, MSP1 shows a steady increase across the full age range with no decline. This is consistent with MSP1 eliciting durable antibody responses (Akpogheneta et al., 2008; Wipasa et al., 2010). The change-point at $\zeta=11.6$ years marks an acceleration in the rate of increase. While the precise mechanism underlying this shift is unclear, plausible explanations include cumulative exposure effects associated with school attendance, immune maturation during early adolescence, or cohort effects reflecting temporal changes in malaria transmission intensity. Distinguishing among these possibilities would require longitudinal data or additional historical information on transmission patterns and intervention coverage.
@@ 7 Discussion
We have introduced a latent variable modelling framework that offers four key methodological contributions. First, it provides a principled way to model the full distribution of antibody concentrations without dichotomization, which remains the dominant practice in serological analyses (Corran et al., 2007; Drakeley et al., 2005; Arnold et al., 2014; Cox et al., 2022) despite well-established information loss (Royston et al., 2006; Kyomuhangi and Giorgi, 2021). By modelling antibody levels directly, the framework provides a more refined characterisation of how immune responses vary and evolve with age. Second, it offers a flexible approach for incorporating age-dependent structure through either mechanistic specifications that encode epidemiological assumptions about exposure and immune response dynamics, or data-driven specifications that adaptively capture empirical patterns without imposing strong structural constraints. Third, by explicitly representing immune activation as a continuum, the framework more naturally accommodates complex population heterogeneity in immunological responses. Fourth, we introduce an $L_{2}$ histogram-based estimator that provides a computationally efficient alternative to maximum likelihood estimation, enabling practical implementation of the framework even in computationally demanding contexts with large sample sizes.
@@
When incorporating age into the distribution of the latent variable $T$, the choice between a single Beta distribution and a finite mixture of Beta distributions should be driven primarily by the inferential objectives of the analysis rather than by the need for increased distributional flexibility per se. Although a mixture offers greater flexibility, the trade-off with identifiability remains a practical concern. More fundamentally, the two approaches may yield near-identical fits while carrying qualitatively different interpretations. For example, a U-shaped distribution on $(0,1)$ can be expressed equivalently using a single Beta with $\alpha<1,\beta<1$ or as a two-component mixture concentrated near the boundaries; however, only the latter allows explicit modelling of subpopulations with distinct seroreactivity profiles whose age-dependent prevalence can be linked to mechanistic formulations such as the catalytic model in (11). The single Beta distribution might be a more natural choice when the goal is a descriptive characterization of how age shapes the distribution of $T$ and, hence, the observed antibody levels. The mixture formulation instead becomes advantageous when the scientific question explicitly concerns disentangling immunologically distinct subgroups within the population, and is most naturally applied in contexts where the age dynamics are sufficiently well understood to support a mechanistic parameterization of the mixing probabilities, as in the malaria case studies considered here.
When using a mixture distribution to define the distribution of the latent variable $T$, an open question remains on how many components one should use. In particular, whether this number should be fixed a priori or determined from the data depends fundamentally on the inferential goal. In applications directed at explaining mechanisms of antibody acquisition, as in Section 3.2, the number of components is a deliberate scientific choice: the two-component specification encodes a biologically interpretable contrast between individuals with low and high immune activation, and its justification comes from subject-matter knowledge rather than data-adaptive criteria. However, when the framework is extended to the joint modelling of responses to multiple antigens—on which we provide more discussion below—the goal may lie in the identification of latent immunological subgroups whose number is not known in advance. In this setting, modelling the multivariate latent state through a mixture distribution offers a conceptually appealing alternative to standard clustering methods, since the clustering is performed in a continuous latent space rather than directly on noisy antibody measurements. For such applications, data-driven selection of the number of components is both meaningful and, we believe, practically feasible. Future research could thus extend the current framework to multivariate latent clustering by adapting existing approaches for selecting the number of mixture components, both in a Bayesian and frequentist context (see for example Richardson and Green 1997; Rousseau and Mengersen 2011; Malsiner-Walli et al. 2016).
@@
An important consideration for applying our framework concerns vaccination status. For diseases conferring durable immunity following vaccination or natural infection — e.g. yellow fever, where a single vaccine dose provides lifelong protection (Staples et al., 2015) — binary serological classification may suffice for most purposes, and continuous antibody modelling may offer limited added value. However, for diseases characterised by partial or waning immunity, such as malaria, continuous antibody measurements provide information about immune activation dynamics that is lost when data are dichotomised. The recent introduction of malaria vaccines (RTS,S/AS01 and R21/Matrix-M) (RTS,S Clinical Trials Partnership, 2015; Datoo et al., 2021) presents new challenges for serological surveillance that affect any analytical approach, not just our modelling framework. Because these vaccines target the pre-erythrocytic circumsporozoite protein rather than the blood-stage antigens we analyse (AMA1, MSP1), vaccination does not directly boost these antibody responses. However, reduced transmission resulting from vaccination programmes could alter natural boosting patterns, thereby affecting the age-dependent dynamics we model. Our simulation study provides evidence on the potential consequences of such departures. When interventions uniformly elevate antibody levels across all ages, the estimated age-dependent structure remains largely intact, with misspecification absorbed primarily into the parameters governing the conditional distribution of antibody levels given the latent seroreactivty state. By contrast, when interventions selectively reduce exposure within specific age groups, inference on the age-dependent profile of seroreactivity becomes unreliable, with biases in the estimated age structure that persist regardless of sample size. In such settings, vaccination status should be incorporated into the model either as a covariate or through a stochastic process affecting the latent immune state $T$, depending on the research question and disease context. When data on multiple antigens are available, joint modelling of their responses might further help to distinguish vaccine-induced from infection-induced seroreactivity.
@@
Future research will focus on extending the current modelling framework to address a wider range of questions in sero epidemiological research, including the development of joint models for multiple antibody responses. One natural direction is to replace the single latent quantity with a multivariate latent structure and to combine this with an appropriate multivariate Beta distribution, for example using the construction proposed by Arnold and Ng (2011). This would allow the model to capture dependencies between antibody responses to different antigens and provide deeper insight into how immune markers coevolve following exposure. The same idea extends naturally beyond serology, since the latent variable formulation is directly relevant to high dimensional omics settings such as proteomics and transcriptomics, where disease states are typically treated as discrete categories, for example healthy versus diseased or sepsis molecular endotypes. Recent work using neural network models for disease classification from multi omics data (Hartman et al., 2023) and large transcriptomic and proteomic studies of sepsis subtypes (Wong et al., 2012; Reyes et al., 2020) all adopt this discrete label perspective, despite the underlying host response being biologically continuous rather than strictly partitioned. Within our framework this continuum can be represented by a latent variable $T$, with each biomarker, here each transcript or protein, modelled as an observed outcome $Y_{j}$ whose conditional distribution depends smoothly on $T$ through marker specific mean and variance functions that describe how expression evolves across the response axis. While a single latent variable provides a parsimonious representation of the dominant gradient of biological variation, many omics settings involve several distinct processes, for example inflammation, immune suppression, and endothelial injury, which may jointly determine the multivariate profile. In such cases it may be necessary to adopt a multivariate latent representation and to define the conditional distribution of $Y_{j}$ as a function of this collection of latent quantities, allowing the framework to capture more complex patterns and multiple dimensions of biological variation that cannot be reduced to a single continuum.
@@ Acknowledgements
Let $f_{\vartheta,n}:=\;f_{n}(\cdot;\vartheta)$. We apply Corollary 3.2.3 of van der Vaart and Wellner [1996]. This corollary establishes consistency provided that: (i) the empirical criterion converges uniformly to the population criterion, i.e., $\sup_{\vartheta\in\vartheta}|\mathbb{M}_{n}(\vartheta)-\mathbb{M}(\vartheta)|\xrightarrow{p}0$. and that (ii) the population maximum is well-separated. Assumption A2. directly gives condition (ii). Thus, it remains only to show condition (i).
We prove this condition in two step. We first establish that the unparameterized histogram $\widehat{f}_{n}$ converges to the true density $f_{0}$ in the $L_{2}$ norm. Freedman and Diaconis [1981] provide the exact mean integrated squared error (MISE) decomposition in their Proposition (1.10):
@@
where $f_{0,n}$ is the theoretical discretized density (the expectation of $\widehat{f}_{n}$). The first term is bounded by $(nh_{n})^{-1}$ and under Assumption A4., this term vanishes. The second term corresponds to the approximation error of the binned density; as shown in Eq. (2.5) of Freedman and Diaconis [1981], this vanishes as $h_{n}\to 0$ for any $f_{0}\in L_{2}$. Therefore, $\mathbb{E}[\|\widehat{f}_{n}-f_{0}\|_{2}^{2}]\to 0$, and by Markov’s inequality:
@@
Again, following Assumption A3. and (16), the terms in parentheses are bounded uniformly and bounded in probability (note that if two terms tend to zero in probability, so does their product). Freedman and Diaconis [1981] show in Proposition (2.7) that when the density is absolutely continuous with square-integrable derivative, then $\|f_{\vartheta,n}-f(\cdot;\vartheta)\|^{2}_{2}$ is $O(h_{n}^{2})$. Specifically, for small $h_{n}$, the term is bounded by a constant times $h_{n}^{2}\|f^{\prime}(\cdot;\vartheta)\|_{2}^{2}$. Thus under Assumption A5., $\sup_{\vartheta\in\vartheta}\|f^{\prime}(\cdot;\vartheta)\|_{2}<\infty$: